IfON=8x−3,LM=7x+6,NM=x−6,and
OL =5y + 4, find the values of x and y given that LMNO is a parallelogram.

Answer by fractalier(6550)

ON is opposite and equal to LM, so
8x - 3 = 7x + 6
x = 9
Then NM = OL and if x = 9
we have
x - 6 = 5y + 4
9 - 6 = 5y + 4
3 = 5y + 4
-1 = 5y
y = -1/5